# Transversal Hamilton cycles in digraph collections

### v1

@@ 5 Extremal case
In this case, each vertex in $Y\setminus(X\cup X_{Y}^{2})$ belongs to $Y_{i}$ for at least $(1-13\sqrt{\delta})|\mathcal{C}_{2}|$ colors $i\in\mathcal{C}_{2}$. Recall that $|B|-|A|=r$ with $0\leq r\leq(8{\delta}+\sqrt{\epsilon})n$. The proof is divided into four steps.
@@
In this step, we claim that by moving vertices in $B\cap V_{\rm bad}$ or deleting rainbow directed paths inside $\mathcal{D}[B]$, we can make the number of remaining vertices in $B$ and $A$ differ by $\sigma$, where $\sigma=1$ if $|A|+|B|$ is odd and $\sigma=0$ otherwise. Denote $s:=|V_{\rm bad}\cap B|$. If $|B|-|A|-\sigma\leq 2s$, then move $\frac{|B|-|A|-\sigma}{2}$ vertices in $V_{\rm bad}\cap B$ to $A$, and our desired result holds. Hence, it suffices to consider $|B|-|A|-\sigma>2s$.
Now, we move all vertices in $V_{\rm bad}\cap B$ to $A$. Hence, $B\cap V_{\rm bad}=\emptyset$. Assume that $\{Q_{1},\ldots,Q_{t}\}$ is a set of disjoint maximal rainbow directed paths in $\mathcal{D}[B]$. If $|E(Q_{1})|+\cdots+|E(Q_{t})|\geq|B|-|A|-\sigma$, then there must exist a set of disjoint rainbow directed paths, say $\{Q_{1}^{\prime},\ldots,Q_{t^{\prime}}^{\prime}\}$, such that $|E(Q_{1}^{\prime})|+\cdots+|E(Q_{t^{\prime}}^{\prime})|=|B|-|A|-\sigma$. That is, $|A|-t^{\prime}=|B|-(|V(Q_{1}^{\prime})|+\cdots+|V(Q_{t^{\prime}}^{\prime})|)-\sigma$. Using Claim 5.3 (ii), we connect $Q_{1}^{\prime},\ldots,Q_{t^{\prime}}^{\prime}$ into a single rainbow directed path $P^{1}$ with length at most $27\delta n$, starting at $A$ and ending at $B$. (In fact, by using Claim 5.3 (ii), we only get a rainbow path with two endpoints in $B$. But we usually want to find a rainbow directed path with endpoints in different parts. Hence one may extend it by using an unused color of $\mathcal{C}_{2}$ and an unused vertex in $A$). Therefore, $|A\setminus V(P^{1})|=|B\setminus V(P^{1})|-\sigma$, as desired.
In what follows, it suffices to consider $|E(Q_{1})|+\cdots+|E(Q_{t})|<|B|-|A|-\sigma=r-2s-\sigma$. Now, we move all vertices of $Q_{1},\ldots,Q_{t}$ to $A$. Let $\mathcal{C}^{\prime}:=\mathcal{C}\setminus{\rm col}(\bigcup_{i\in[t]}Q_{i})$. Hence
@@
By the maximality of $\{Q_{1},\ldots,Q_{t}\}$, one has $D_{i}[B]=\emptyset$ for all $i\in\mathcal{C}^{\prime}$. It follows from $\delta^{0}(\mathcal{D})\geq\left\lceil\frac{n}{2}\right\rceil$ that $\left\lceil\frac{n}{2}\right\rceil\leq|A|+|X^{\prime}|\leq\frac{n}{2}+19{\delta}n$. Now, by Lemma 5.14 (which is presented in the final of this subsection), we obtain that $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
@@
Step 2. Construct a series of disjoint rainbow directed rainbow $P_{3}$ such that all centers of them are exactly all vertices in $V_{\rm bad}\cup X^{\prime}$.
Recall that $|X_{A}^{2}\cup X_{B}^{2}|\leq|X_{A}\cup X_{B}|<\frac{1}{2}\sqrt{\delta}n$. Let $v\in X_{Y}^{2}$. If $v\in Z$, then $v$ is moved from $B$ to $A$ in Step 1, so $v\in X_{B}$. Therefore, $v\in A_{i}\cup B_{i}$ for at least $(1-6\sqrt{\delta})|\mathcal{C}_{2}|$ colors $i\in\mathcal{C}_{2}$ and $v\not\in B_{i}$ for at least $10\sqrt{\delta}|\mathcal{C}_{2}|$ colors $i\in\mathcal{C}_{2}$. Hence, $v\in A_{i}$ for at least $4\sqrt{\delta}|\mathcal{C}_{2}|-27\delta n>4|V_{\rm bad}\cup X^{\prime}|$ colors $i\in\mathcal{C}_{2}\setminus{\rm col}(P^{1})$. If $v\in Y$, then $v$ belongs to $Y_{i}$ for at least $(3\sqrt{\delta}-27\delta)n>4|V_{\rm bad}\cup X^{\prime}|$ colors $i\in\mathcal{C}_{2}\setminus{\rm col}(P^{1})$. Hence, using colors in $\mathcal{C}_{2}\setminus{\rm col}(P^{1})$, we can greedily choose $|X_{A}^{2}\cup X_{B}^{2}|$ disjoint rainbow directed $P_{3}$, such that their centers correspond to all vertices in $X_{A}^{2}\cup X_{B}^{2}$. Moreover, if $v\in(X_{A}^{2}\cup X_{B}^{2})\cap Y$, then the rainbow directed $P_{3}$ with center $v$ has endpoints in $Z\setminus(V_{\rm bad}\cup X^{\prime}\cup V(P^{1}))$.
For vertices in $X^{\prime}$, we know from its definition that by avoiding vertices and colors in $P^{1}$ and all rainbow directed paths chosen for vertices in $X_{A}^{2}\cup X_{B}^{2}$, there exists a set of disjoint rainbow directed $P_{3}$ with centers in $X^{\prime}$ and endpoints in different parts.
By using Claim 5.3 (i)-(ii), we connect $P^{1}$, all rainbow directed $P_{3}$ with centers in $X_{A}^{2}\cup X_{B}^{2}\cup X^{\prime}$ (as obtained above) into a single rainbow directed path. This results in a rainbow directed path $P^{2}=u_{1}\ldots v_{2}$ with $v_{2}\in B\setminus V_{\rm bad}$ and length at most

### v2

@@ 1. $(\dagger)$
To prove Theorem 5.2, we need the following lemma. This establishes that every digraph collection $\mathcal{D}$ contains a transversal directed Hamilton cycle if it admits an almost balanced partition, with the small parts inducing an empty graph for almost all colors. We first state a preliminary result from [11].
Lemma 6.4 ([11]).
Let $\mathcal{C}$ be a set of colors, and $\mathcal{G}=\{G_{i}[Y,B]:i\in\mathcal{C}\}$ be a collection of bipartite graphs with the common bipartition $Y\cup B$ such that $7|Y|<|B|\leq\frac{3}{5}|\mathcal{C}|$. If $\sum_{i\in\mathcal{C}}|E(D_{i}[Y,B])|\geq t|B||\mathcal{C}|$ for some integer $t$ with $1\leq t\leq|Y|$, then $\mathcal{G}$ contains $t$ disjoint rainbow stars on $5$ vertices, and each of them has its center in $Y$ and other vertices in $B$.
Lemma 6.5.
Proof.
$=(|A|-\delta^{\frac{1}{2}}|Y_{1}|)|B||\mathcal{C}^{\prime}|.$
It implies that $|Y_{1}|\leq\frac{\gamma n}{\sqrt{\delta}}\leq 3\sqrt{\delta}n$. Similarly, $|Y_{2}|\leq\frac{\gamma n}{\sqrt{\delta}}\leq 3\sqrt{\delta}n$. For a vertex $v\in Y_{1}\setminus Y_{2}$ (resp. $v\in Y_{2}\setminus Y_{1}$), we have $d_{i}^{-}(v,B)>(1-\delta^{\frac{1}{4}})|B|$ (resp. $d_{i}^{+}(v,B)>(1-\delta^{\frac{1}{4}})|B|$) for at least $(1-\delta^{\frac{1}{4}})|\mathcal{C}^{\prime}|$ colors $i\in\mathcal{C}^{\prime}$. Notice that Lemma 6.2 and Lemma 6.3 (i)-(ii) hold by setting $V_{\rm bad}:=Y_{1}\cup Y_{2}$ and $\mathcal{C}_{2}=\mathcal{C}^{\prime}$.
Case 1. $|Y_{1}|,|Y_{2}|\geq\gamma n$.
In view of Lemma 6.4, there exists a subset $Y_{1}^{\prime}\subseteq Y_{1}$ with size $\gamma n$ such that vertices in $Y_{1}\setminus Y_{1}^{\prime}$ can be covered by disjoint rainbow $P_{3}$ copies inside $\{D_{i}[Y_{1},B]:i\in\mathcal{C}^{\prime}\}$ with centers in $Y_{1}\setminus Y_{1}^{\prime}$; and a subset $Y_{2}^{\prime}\subseteq Y_{2}$ with size $\gamma n$ such that vertices in $Y_{2}\setminus Y_{2}^{\prime}$ can be covered by disjoint rainbow $P_{3}$ copies inside $\{D_{i}[B,Y_{2}]:i\in\mathcal{C}^{\prime}\}$ with centers in $Y_{2}\setminus Y_{2}^{\prime}$. Recall that each vertex in $Y_{1}$ (resp. $Y_{2}$) has at least $\left\lceil\frac{n}{2}\right\rceil-(1-\delta^{\frac{1}{4}})|B|>4|Y|$ out-neighbors (resp. in-neighbors) in $A$ for at least $\delta^{\frac{1}{4}}|\mathcal{C}^{\prime}|>\delta^{\frac{1}{4}}(1-\delta)n>4|Y|$ digraphs $D_{i}$ with $i\in\mathcal{C}^{\prime}$. Therefore, by using colors in $\mathcal{C}^{\prime}$,
  • vertices in $(Y_{1}\cup Y_{2})\setminus(Y_{1}^{\prime}\cup Y_{2}^{\prime})$ can be covered by disjoint rainbow directed $P_{3}$ copies with centers in $(Y_{1}\cup Y_{2})\setminus(Y_{1}^{\prime}\cup Y_{2}^{\prime})$ and endpoints in $B$,
  • vertices in $Y_{1}^{\prime}\setminus Y_{2}^{\prime}$ can be covered by disjoint rainbow directed $P_{3}$ copies starting at $B$ and ending at $A$ with centers in $Y_{1}^{\prime}\setminus Y_{2}^{\prime}$,
  • vertices in $Y_{2}^{\prime}\setminus Y_{1}^{\prime}$ can be covered by disjoint rainbow directed $P_{3}$ copies starting at $A$ and ending at $B$ with centers in $Y_{2}^{\prime}\setminus Y_{1}^{\prime}$,
  • vertices in $Y_{1}^{\prime}\cap Y_{2}^{\prime}$ can be covered by disjoint rainbow directed $P_{3}$ copies with centers in $Y_{1}^{\prime}\cap Y_{2}^{\prime}$ and endpoints in $A$.
For each $y\in Y_{1}\cup Y_{2}$, let $P_{y}:=y^{1}yy^{2}$ denote the rainbow directed $P_{3}$ with center $y$, where $P_{y}$ has colors $c_{y}^{1}$ and $c_{y}^{2}$. Let $\mathbf{P}=\{P_{y}:y\in Y_{1}\cup Y_{2}\}$. By using Lemma 6.3, one may connect all of the above rainbow directed $P_{3}$ copies into a single rainbow directed path $P^{1}$ that starts at $A$ and ends at $B$. Then
It is routine to check that there exists a rainbow matching inside $\{D_{i}^{\pm}[A\setminus V(P^{1}),B\setminus V(P^{1})]:i\in\mathcal{C}^{\prime\prime}\}$, say $M$, such that $D_{j}[A\setminus V(P^{1}\cup M),B\setminus V(P^{1}\cup M)]\neq\emptyset$ for each $j\in{\rm col}(M)$. Furthermore, $D_{j}$ is $(19\sqrt{\delta},{\rm EC}1)$-extremal for all $j\in\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)$.
We greedily choose two disjoint rainbow directed paths $P_{A}$ and $P_{B}$ inside $\{D_{i}[A\setminus V(P^{1}\cup M)]\cup D_{i}[B\setminus V(P^{1}\cup M)]:i\in\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)\}$ with lengths $\lceil\frac{|\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)|}{2}\rceil$ and $\lfloor\frac{|\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)|}{2}\rfloor$ respectively, such that $V(P_{A})\subseteq A$ and $V(P_{B})\subseteq B$. Applying Lemma 6.3 (i), we successively connect $P^{1}$, all rainbow edges in $M$, $P_{A}$ and $P_{B}$ into a single rainbow path $P^{2}$ whose endpoints lie in different parts. Clearly, $|A\setminus V(P^{2})|=|B\setminus V(P^{2})|$. Together with Lemma 6.2, we know that $\mathcal{D}$ contains a transversal directed Hamilton cycle, as desired.
Subcase 1.2. $n$ and $|\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)|$ have different parity.
Notice that $|B|=\lfloor\frac{n}{2}\rfloor-\gamma n$. Hence for each $D_{i}$ with $i\in\mathcal{C}$, each vertex in $B$ has at least $\gamma n+1$ in-neighbors in $A$. Therefore, each vertex in $B$ has at least one in-neighbor outside $Y_{1}^{\prime}$.
We first assume $|\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)|\geq 1$. Choose an unused vertex $y\in B$ and a color $c\in\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)$. Assume that ${xy}\in E(D_{c}[A,B])$ with $x\notin Y_{2}^{\prime}$. If $x\in A\setminus(Y_{1}\cup Y_{2})$, then $x$ can be chosen outside $V(P^{1}\cup M)$. Using Lemma 6.3 (i)-(ii), we connect $xy$ and $P^{1}$ into a single rainbow directed path $P$. If $x\in(Y_{1}\cup Y_{2})\setminus Y_{1}^{\prime}$, then we replace the end vertex of $P_{x}$ by $y$. In either situation the problem reduces to Subcase 1.1, yielding a transversal directed Hamilton cycle in $\mathcal{D}$, as desired.
Now, we consider that $n$ is odd and $|\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)|=0$. For each $D_{i}$ with $i\in\mathcal{C}$, each vertex in $A$ has at least one in-neighbor outside $Y_{1}^{\prime}$. Choose an unused vertex $y\in A\setminus(Y_{1}\cup Y_{2})$ and a color $c\in\mathcal{C}$. Then there exists a vertex $x\in A\setminus Y_{1}^{\prime}$ such that ${xy}\in E(D_{c}[A])$. By a similar discussion as above, we obtain a transversal directed Hamilton cycle inside $\mathcal{D}$, as desired.
Case 2. At least one of $Y_{1}$ and $Y_{2}$ has size less than $\gamma n$.
In this case, we consider only $|Y_{1}|\geq\gamma n$ and $|Y_{2}|<\gamma n$; the remaining two cases are analogous. Following the argument of Case 1, we know that all vertices in $Y_{1}$ can be covered by a set of disjoint rainbow directed copies of $P_{3}$ with centers in $Y_{1}$ and endpoints in $A\setminus(Y_{1}\cup Y_{2})$. Furthermore, each vertex in $Y_{2}\setminus Y_{1}$ can be covered by a rainbow directed $P_{3}$ with its center in $Y_{2}\setminus Y_{1}$, starting from $A\setminus(Y_{1}\cup Y_{2})$ and using colors in $\mathcal{C}^{\prime}$. (In this process, for the choice of the other endpoint of each rainbow directed $P_{3}$, we prefer using vertices in $A$ over vertices in $B$.) By using unused colors in $\mathcal{C}$ and unused vertices in $A\setminus Y$, one may extend those rainbow directed $P_{3}$ as long as possible or choose all other disjoint rainbow directed paths. Let $\mathbf{P}=\{Q_{1},Q_{2},\ldots,Q_{t}\}$ be a set consisting of all disjoint rainbow directed paths in the above, where each $Q_{i}$ has at least one endpoint in $A\setminus Y$ and length $s_{i}\ (1\leq i\leq t)$ inside $A$. We next consider the following two possible cases.
Subcase 2.1. $|B|\geq|A|-(s_{1}+\cdots+s_{t})-\sigma$.
In this subcase, there exists a set $\mathbf{P}^{\prime}=\{Q_{1}^{\prime},Q_{2}^{\prime},\ldots,Q_{\ell}^{\prime}\}$ such that $|B|=|A|-(s_{1}^{\prime}+\cdots+s_{\ell}^{\prime})-\sigma$, the endpoints of each $Q_{i}^{\prime}$ lie outside $Y$ and $Y_{2}\cup Y_{1}^{\prime}\subseteq V(Q_{1}^{\prime}\cup\ldots\cup Q_{\ell}^{\prime})$, where $s_{i}^{\prime}=|E(Q_{i}^{\prime}[A])|$ for each $i\in[\ell]$. By Lemma 6.3 (i)-(ii), we connect all rainbow directed paths in $\mathbf{P}^{\prime}$ into a single rainbow directed path $P^{1}$, whose endpoints are in different parts. Since $|A|-|B|=2\gamma n+\sigma$, we have $\ell\leq s_{1}^{\prime}+\cdots+s_{\ell}^{\prime}=2\gamma n$. Therefore, $|E(P^{1})|\leq 8\gamma n$ and $|A\setminus V(P^{1})|-\sigma=|B\setminus V(P^{1})|$.
Recall that all vertices in $Y_{1}\setminus V(\mathbf{P}^{\prime})$ can also be covered by a set of disjoint rainbow directed copies of $P_{3}$ with centers in $Y_{1}\setminus V(\mathbf{P}^{\prime})$ and both endpoints in $B$. Applying Lemma 6.3 (i) again to connect those rainbow directed paths with $P^{1}$, we get a rainbow directed path $P^{2}$ with length at most $8\gamma n+12\sqrt{\delta}n$, and with endpoints in different parts.
Choose a maximum rainbow matching inside $\{D_{i}^{\pm}[A\setminus V(P^{2}),B\setminus V(P^{2})]:i\in\mathcal{C}^{\prime\prime}\setminus{\rm col}(P^{2})\}$, say $M$, such that $D_{j}^{\pm}[A\setminus V(P^{2}\cup M),B\setminus V(P^{2}\cup M)]\neq\emptyset$ for each $j\in{\rm col}(M)$. Clearly, $D_{j}$ is $(13\sqrt{\delta},{\rm EC}1)$-extremal for all $j\in\mathcal{C}^{\prime\prime}\setminus{\rm col}(P^{1}\cup M)$. In fact, when constructing $P^{1}-B-Y$, we may prioritize the use of colors from $\mathcal{C}^{\prime\prime}\setminus{\rm col}(P^{1}\cup M)$ over other colors. Hence either ${\rm col}(P^{1}-B-Y)\subseteq\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)$ or $\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)\subseteq{\rm col}(P^{1}-B-Y)$.
Based on Lemma 6.3 (i), one may connect $P^{1}$ and all edges of $M$ into a single rainbow directed path $P^{2}$, whose endpoints are in different parts and length is at most $8\gamma n+4{\delta}n$. Next, we are to choose two disjoint rainbow paths $P_{A}$ and $P_{B}$ inside $\{D_{i}[A\setminus V(P^{1}\cup M)]\cup D_{i}[B\setminus V(P^{1}\cup M)]:i\in\mathcal{C}^{\prime\prime}\setminus{\rm col}(P^{1}\cup M)\}$ such that $V(P_{A})\subseteq A$ and $V(P_{B})\subseteq B$ respectively, whose lengths are determined by the parity of $|\mathcal{C}^{\prime\prime}\setminus{\rm col}(P^{1}\cup M)|$. Clearly, if $n$ and $|\mathcal{C}^{\prime\prime}\setminus{\rm col}(P^{1}\cup M)|$ have the same parity, then by Lemma 6.2, we know that $\mathcal{D}$ contains a transversal directed Hamilton cycle, as desired. Hence $n$ and $|\mathcal{C}^{\prime\prime}\setminus{\rm col}(P^{1}\cup M)|$ have different parity. By an argument analogous to that in the proof of Step 4 in Case 1, we can obtain the desired transversal Hamilton cycle in $\mathcal{D}$.
Subcase 2.2. $|B|<|A|-(s_{1}+\cdots+s_{t})-\sigma$.
Now, we are ready to give the proof of Theorem 5.2.
Proof of Theorem 5.2.
Choose constant $\epsilon,\delta,\eta$ such that
$\frac{1}{n}\ll\epsilon\leq\delta^{2}\ll\eta\ll 1,$
let $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ be a collection of digraphs on a common vertex set $V$ of size $n$ and $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$. Suppose that $\mathcal{D}$ does not contain transversal directed Hamilton cycles. By Lemma 5.1, for each $i\in[m]$, $D_{i}$ has a characteristic partition $({A_{i}},{B_{i}},{L_{i}})$ satisfying (C1)-(C4). Moreover, for every $i\in[m]$ and $Y\in\{A,B\}$, we have
$|L_{i}|\leq 2\epsilon n\ \text{and}\ |{Y}_{1}{\triangle}Y_{i}|<2\delta n.$ (2)
Expand $A_{1}\cup B_{1}$ to an equitable partition $A\cup B$ of $V$. For simplicity, we assume $Y\in\{A,B\}$ and $Z\in\{A,B\}\setminus\{Y\}$ in the following. Let $\mathcal{C}_{1}\cup\mathcal{C}_{2}$ be a partition of $[m]$ with
$\mathcal{C}_{1}=\{i\in[m]:D_{i}\penalty 10000\ \textrm{is}\penalty 10000\ (\epsilon,{\rm EC}1){\text{-extremal}}\},$
$\mathcal{C}_{2}=\{i\in[m]:D_{i}\penalty 10000\ \textrm{is}\penalty 10000\ (\epsilon,{\rm EC}2){\text{-extremal}}\}.$
Let $\hat{\mathcal{C}}:=\bigcup_{k\in[2]}\psi(\mathcal{C}_{k}),$ where $\psi(\mathcal{C}_{k})=\mathcal{C}_{k}$ if $|\mathcal{C}_{k}|\geq\eta n$ and $\psi(\mathcal{C}_{k})=\emptyset$ otherwise. Clearly, $|\hat{\mathcal{C}}|\geq(1-4\sqrt{\delta}-\eta)n$. Define
$X:=$ $\{x\in V:x\not\in A_{i}\cup B_{i}\penalty 10000\ \textrm{for at least}\penalty 10000\ 6\sqrt{\delta}|\hat{\mathcal{C}}|\penalty 10000\ \textrm{colors}\penalty 10000\ i\in\hat{\mathcal{C}}\}$,
$X_{Y}:=$ $\{x\in Y\setminus X:x\not\in Y_{i}\penalty 10000\ \textrm{for at least}\penalty 10000\ 10\sqrt{\delta}|\hat{\mathcal{C}}|\penalty 10000\ \textrm{colors}\penalty 10000\ i\in\hat{\mathcal{C}}\}$.
[…]
$X_{Y}^{k}:=\left\{\begin{array}[]{ll}\{x\in X_{Y}:x\in Y_{i}\penalty 10000\ \textrm{for at least}\penalty 10000\ 3\sqrt{\delta}n\penalty 10000\ \textrm{colors}\penalty 10000\ i\in{\mathcal{C}_{k}}\},&\textrm{if $|\mathcal{C}_{k}|\geq\eta n$,}\\[5.0pt] \emptyset,&\textrm{otherwise.}\end{array}\right$.
[…]
It follows that $|X_{Y}\setminus(\bigcup_{k\in[2]}X_{Y}^{k})|<4\delta n$. Then we move vertices in $X_{A}\setminus(\bigcup_{k\in[2]}X_{A}^{k})$ to $B$ and vertices in $X_{B}\setminus(\bigcup_{k\in[2]}X_{B}^{k})$ to $A$.
[…]
@@
In this case, each vertex in $Y\setminus(X\cup X_{Y}^{2})$ belongs to $Y_{i}$ for at least $(1-13\sqrt{\delta})|\mathcal{C}_{2}|$ colors $i\in\mathcal{C}_{2}$. Recall that |B||A|=r|B|-|A|=r with $0\leq r\leq(8{\delta}+\sqrt{\epsilon})n$. The proof is divided into four steps.
@@
In this step, we claim that by moving vertices in $B\cap V_{\rm bad}$ or deleting rainbow directed paths inside $\mathcal{D}[B]$, we can make the number of remaining vertices in $B$ and $A$ differ by $\sigma$, where $\sigma=1$ if $|A|+|B|$ is odd and $\sigma=0$ otherwise. Denote $s:=|V_{\rm bad}\cap B|$. If $|B|-|A|-\sigma\leq 2s$, then move $\frac{|B|-|A|-\sigma}{2}$ vertices in $V_{\rm bad}\cap B$ to $A$, and our desired result holds. Hence, it suffices to consider $|B|-|A|-\sigma>2s$. Now, we move all vertices in $V_{\rm bad}\cap B$ to $A$. Hence, $B\cap V_{\rm bad}=\emptyset$. Assume that $\{Q_{1},\ldots,Q_{t}\}$ is a set of disjoint maximal rainbow directed paths in $\mathcal{D}[B]$.
On the one hand, suppose that $|E(Q_{1})|+\cdots+|E(Q_{t})|\geq|B|-|A|-\sigma$. Then there must exist a set of disjoint rainbow directed paths, say $\{Q_{1}^{\prime},\ldots,Q_{t^{\prime}}^{\prime}\}$, such that $|E(Q_{1}^{\prime})|+\cdots+|E(Q_{t^{\prime}}^{\prime})|=|B|-|A|-\sigma$. That is, $|A|-t^{\prime}=|B|-(|V(Q_{1}^{\prime})|+\cdots+|V(Q_{t^{\prime}}^{\prime})|)-\sigma$. Using Lemma 6.3 (ii), we connect $Q_{1}^{\prime},\ldots,Q_{t^{\prime}}^{\prime}$ into a single rainbow directed path $P^{1}$ with length at most $27\delta n$, starting at $A$ and ending at $B$. (In fact, by using Lemma 6.3 (ii), we only get a rainbow path with two endpoints in $B$. But we usually want to find a rainbow directed path with endpoints in different parts. Hence one may extend it by using an unused color of $\mathcal{C}_{2}$ and an unused vertex in $A$). Therefore, $|A\setminus V(P^{1})|=|B\setminus V(P^{1})|-\sigma$, as desired.
On the other hand, suppose that $|E(Q_{1})|+\cdots+|E(Q_{t})|<|B|-|A|-\sigma=r-2s-\sigma$. In this case, we move all vertices of $Q_{1},\ldots,Q_{t}$ to $A$. Let $\mathcal{C}^{\prime}:=\mathcal{C}\setminus{\rm col}(\bigcup_{i\in[t]}Q_{i})$. Hence
@@
By the maximality of $\{Q_{1},\ldots,Q_{t}\}$, one has $D_{i}[B]=\emptyset$ for all $i\in\mathcal{C}^{\prime}$. It follows from $\delta^{0}(\mathcal{D})\geq\left\lceil\frac{n}{2}\right\rceil$ that $\left\lceil\frac{n}{2}\right\rceil\leq|A|+|X^{\prime}|\leq\frac{n}{2}+19{\delta}n$. Now, by Lemma 6.5, we obtain that $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
@@
Step 2. Construct a sequence of disjoint rainbow directed $P_{3}$ copies such that all centers of them are exactly all vertices in $V_{\rm bad}\cup X^{\prime}$.
Recall that $|X_{A}^{2}\cup X_{B}^{2}|\leq|X_{A}\cup X_{B}|<\frac{1}{2}\sqrt{\delta}n$. Let $v\in X_{Y}^{2}$. If $v\in Z$, then $v$ is moved from $B$ to $A$ in Step 1, so $v\in X_{B}$. Therefore, $v\in A_{i}\cup B_{i}$ for at least $(1-6\sqrt{\delta})|\mathcal{C}_{2}|$ colors $i\in\mathcal{C}_{2}$ and $v\not\in B_{i}$ for at least $10\sqrt{\delta}|\mathcal{C}_{2}|$ colors $i\in\mathcal{C}_{2}$. Hence, $v\in A_{i}$ for at least $4\sqrt{\delta}|\mathcal{C}_{2}|-27\delta n>4|V_{\rm bad}\cup X^{\prime}|$ colors $i\in\mathcal{C}_{2}\setminus{\rm col}(P^{1})$. If $v\in Y$, then $v$ belongs to $Y_{i}$ for at least $(3\sqrt{\delta}-27\delta)n>4|V_{\rm bad}\cup X^{\prime}|$ colors $i\in\mathcal{C}_{2}\setminus{\rm col}(P^{1})$. Hence, using colors in $\mathcal{C}_{2}\setminus{\rm col}(P^{1})$, we can greedily choose $|X_{A}^{2}\cup X_{B}^{2}|$ disjoint rainbow directed $P_{3}$ copies, such that their centers correspond to all vertices in $X_{A}^{2}\cup X_{B}^{2}$. Moreover, if $v\in(X_{A}^{2}\cup X_{B}^{2})\cap Y$, then the rainbow directed $P_{3}$ with center $v$ has endpoints in $Z\setminus(V_{\rm bad}\cup X^{\prime}\cup V(P^{1}))$.
For vertices in $X^{\prime}$, we know from its definition that by avoiding vertices and colors in $P^{1}$ and all rainbow directed paths chosen for vertices in $X_{A}^{2}\cup X_{B}^{2}$, there exists a set of disjoint rainbow directed $P_{3}$ copies with centers in $X^{\prime}$ and endpoints in different parts.
By using Lemma 6.3 (i)-(ii), we connect $P^{1}$, all rainbow directed $P_{3}$ copies with centers in $X_{A}^{2}\cup X_{B}^{2}\cup X^{\prime}$ (as obtained above) into a single rainbow directed path. This results in a rainbow directed path $P^{2}=u_{1}\ldots v_{2}$ with $v_{2}\in B\setminus V_{\rm bad}$ and length at most